### v6

# Short vs. Long-term Coordination of Drones: When Distributed Optimization Meets Deep Reinforcement Learning

Abstract
Swarms of autonomous interactive drones can provide compelling sensing capabilities in Smart City applications, such as traffic monitoring. This paper focuses on the task assignment problem for large-scale spatio-temporal sensing by a drone swarm. However, existing approaches have distinct challenges: evolutionary distributed optimization, such as collective learning, lacks long-term adaptability in dynamic environments, while deep reinforcement learning (DRL) is limited to scale effectively due to the curse of dimensionality. Therefore, this paper proposes a novel synergetic optimization approach by integrating long-term DRL and short-term collective learning. Through this approach, each drone independently and proactively determines its flying direction and recharging location using DRL, while evolving their navigation and sensing policies through collective learning based on a structured tree communication model. Extensive experiments with datasets generated from realistic urban mobility demonstrate an outstanding performance of the proposed solution in complex scenarios. New insights show that this approach provides a win-win synthesis of short-term and long-term strategies for drone-based traffic monitoring, with short-term methods addressing training complexity and energy management, while long-term methods preserving high sensing performance.
Index Terms:
drones, UAV, navigation, sensing, recharging, distributed optimization, deep reinforcement learning
I Introduction
Unmanned Aerial Vehicles (UAVs), referred to as drones, can organize themselves into swarms, fostering collaboration and efficiency in sensor data collection within Smart Cities [1]. With their mobility, autonomy, and diverse sensors, drones have been widely used in several applications, such as monitoring traffic congestion [2] and mapping natural disasters [3]. Despite these advantages, drones are constrained by limited battery capacity, which impacts their spatio-temporal coverage. This limitation necessitates sophisticated autonomous control systems and precise task assignment within drone swarms. This involves the assignment of different sensing tasks to each drone while meeting the sensing requirements, drone capabilities and constraints [4]. Recently, scholars have solved the NP-hard combinatorial optimization problem of task assignment for spatio-temporal sensing, using evolutionary algorithms ranging from genetic algorithms to swarm intelligence methods [4, 5]. To enhance robustness and minimize the effects of individual drone failures, distributed optimization approaches have been introduced. These approaches allow drones to autonomously self-organize and self-assign tasks in a decentralized manner, preserving the autonomy of individual drones. However, these distributed approaches are reactive to the current state of the environment rather than proactive to address long-term operational challenges strategically.
Strategic foresight in navigation and sensing is crucial for optimizing drone operations in dynamic environments, significantly influenced by flying directions and future sensor data requirements. For example, drones aware of an expected increase in traffic can proactively fly to those areas, improving tasks such as vehicle detection, even if the short-term benefits are not evident. This "slower is faster" effect emphasizes the importance of long-term planning for overall effectiveness in drone sensing. To address long-term optimization, deep reinforcement learning (DRL) has been introduced, utilizing the Bellman equation to consider cumulative rewards [6]. DRL based on deep neural networks support drones, as heterogeneous agents, to manage complex state spaces and adapt to changing environments.
However, traditional DRL approaches cannot scale well with a large number of agents and extended time spans due to the curse of dimensionality. With more agents, coordinating them becomes exponentially more complex, leading to computational challenges. Furthermore, DRL methods need to account of environmental uncertainties and the long-term consequences of agents’ actions, which become increasingly complex as the time horizon extends. For example, drones often make flight path decisions over short intervals, such as seconds or minutes, but face significant challenges when required to adapt their policies for a long duration (e.g., a full day). Meanwhile, the extended operation is longer than the maximum flight time of drones, enabling them to return to a base station for recharging and depart from that place for the subsequent task. The selection of these recharging places constraints the drones’ flight ranges, which in turn affects their sensing performance.
Therefore, in this paper, we address a novel task assignment problem for large-scale spatio-temporal sensing, focusing on the navigation (determining flight path), sensing (scheduling data collection) and recharging (choosing charging stations as destinations) of drone swarms in changing environments. To jointly manage these tasks, we propose a new synergetic optimization approach to coordinate drone swarms through Distributed Optimization and (deep) Reinforcement Learning (DO-RL), which involves a multi-agent DRL algorithm that builds upon on an iterative collective learning approach [7, 8]. It allows drones to determine long-term flying directions from departures to destinations (i.e., charging stations) via DRL, while autonomously evolving their short-term navigation and sensing policies based on shared plans from other drones through collective learning. The distributed sensing results generated during operations serve as rewards and observations for further DRL training. These results are aggregated and shared using a structured tree communication model, where each drone interacts with its parent and child nodes, making independent decisions. This ensures scalability (supporting a large number of agents), efficiency (low communication and computational cost), decentralization and resilience [9, 10, 11]. By employing this collaborative approach, collective learning reduces the frequency of decision-making and mitigates inefficient actions, effectively compensating for the limitations of DRL while retaining its strengths in long-term planning and optimization. In addition, a set of auxiliary approaches is introduced, including plan generation, plan selection, periodic state updates and a centralized training and decentralized execution module. To validate DO-RL, this paper conducts experimental evaluations with real-world data and emulated traffic monitoring scenarios, comparing it against state-of-the-art baseline methods.
The contributions of this paper are outlined as follows: (i) The first study of task assignment for large-scale, dynamic spatio-temporal sensing by a swarm of drones, concentrating on optimizing the flight paths, data collection strategies, and recharging. (ii) A novel synergetic optimization approach, DO-RL, that tackles this problem by integrating both multi-agent DRL and distributed optimization based on a structured tree communication model. (iii) A testbed for extensive experimentation with realistic traffic patterns and real-world transport networks. (iv) The validation for the superior performance of the proposed approach over three state-of–the-art methods. (v) An open-source implementation of the proposed approach and an open dataset [12] containing all plans of the studied scenario. They can be used as benchmarks to encourage further research on this problem.
TABLE I: Comparison to related work: criteria covered (✓) or not (✕).
Criteria   Ref.:
[13]
[14]
[15]
[16]
[17]
[18]
[19]
This paper
Decentralization
Long-term efficiency
Scalability
Resilience
Adaptability
II Related Work
Sensing task assignment. The UAV task assignment problem for spatio-temporal sensing has been traditionally defined as the Traveling Salesmen Problem [1, 20]. This involves finding the optimal assignment of tasks to drones at specific locations, while considering constraints such as task urgency, time scheduling, and flying costs. The evolutionary optimization methods, such as particle swarm optimization [21], and genetic algorithm [22], have been introduced to formulate the problem of reaching optimal sensing efficiency in terms of coverage, events detection rate and the cost of flight paths. They rely on adaptive exploration of the solution space to optimize task allocation based on shared information and local interactions. However, these methods are centralized, wherein the failure of a central control station can lead to several system disruptions without autonomous recovery mechanisms [23].
Distributed optimization. Several distributed task assignment algorithms have been proposed. One such algorithm is the robust decentralized task assignment (RDTA) [13] that enhances robustness and reduces communication costs by employing decentralized planning for drones. Another one is the consensus-based bundle algorithm (CBBA) [14], which combines market-based mechanisms and situation awareness to converge and avoid task conflicts among drones. Moreover, a planning-based coordination method using collective learning, named Economic Planning and Optimized Selections (EPOS), enables drones to optimize their navigation and sensing options under battery constraints [15, 24]. However, long-term sensing scenarios pose challenges. Drones tend to prioritize immediate maximum returns in the short term, often neglecting the need for strategic decision-making for overall benefits.
DRL-based drone sensing. The deep reinforcement learning (DRL) algorithm proves to be effective for addressing distributed sensing with drones, managing complex optimization goals, energy constraints, and numerous heterogeneous agents. Ding et al. [16] integrated mobile crowdsensing (i.e., human equipped with mobile sensing devices) into UAV sensing, overcoming the limitations of both human mobility and drone battery life. Omoniwa et al. [25] presented a decentralized DRL approach that improves drone energy efficiency in providing wireless connectivity by sharing information with neighbouring drones. Zhao et al. [17] proposed a comprehensive solution for sustainable urban-scale sensing in open parking spaces. Their solution incorporates the task selection and scheduling using DRL, and adaptive recharging assignment. However, these approaches lack autonomy for drones in self-assignment tasks, reducing adaptability to the real-time environmental changes and increasing vulnerability to agent failures [23]. Additionally, they are confined to a small-scale task assignment problem due to the training complexity.
Evolutionary algorithm and DRL. In this optimization process, the related synergetic optimization algorithms integrate the full learning and optimization capabilities of evolutionary algorithms and DRL. They independently solve sub-problems to generate partial solutions, which are then combined to form a complete solution [26]. Earlier work [27, 28] has proposed optimizing team policies using evolutionary algorithms with team rewards, while simultaneously optimizing individual policies through DRL with agent rewards, thus bypassing the non-stationary problem in multi-agent DRL. Li et al. [18] introduce a hybrid framework using shared observation encoders and independent linear policy representations to enhance collaboration on complex task, ensuring stable evolution through agent-level crossover and mutation. While benchmarks, such as MuJoCo [29] and Starcraft II [30], are useful for evaluation of these approaches, these benchmarks may struggle to adapt to the continuously evolving and unanticipated dynamics typical of real-world scenarios, such as fluctuating traffic patterns in drone-based monitoring systems.
Transportation application. In the application of transportation system, drones rely on accurate vehicle observation to detect early signs of traffic congestion. This enables traffic operators to apply mitigation measures, reducing the carbon footprint in one of the world’s highest-emission sectors. Elloumi et al[31] proposed a road traffic monitoring system that generates adaptive drones trajectories by extracting information about vehicles. Samir et al. [19] leveraged a DRL algorithm to optimize the trajectories of UAVs, with the objective of minimizing the age of information of collected data. While these approaches demonstrate high performance in the detection of traffic vehicles, scalability becomes a challenge since they only deal with a small number of drones within a short duration.
In summary, previous approaches have either neglected the long-term advantages of strategic sensing, or failed to effectively adapt to failures and scale without encountering centralized bottlenecks. To overcome these shortcomings, the proposed approach empowers drones to make informed and proactive decisions based on predicted traffic flow, while enabling them to independently adjust to changing conditions and unexpected events, ensuring both scalability and resilience without relying on centralized control.
III System Model
In this section, we define key concepts of scenarios and then formulate the main performance metrics.
III-A Scenario and Assumptions
Sensing map. Consider a swarm of drones $\mathcal{U}\triangleq\{1,2,...,U\}$ performing sensing missions, such as monitoring vehicles, over a grid that represents a 2D map. In this scenario, a set of grid cells (or points of interest) $\mathcal{N}\triangleq\{1,2,...,N\}$ are uniformly arranged to cover the map. The primary goal of the drones is to coordinate their visits to these cells to collect the required data. Furthermore, a set of charging stations $\mathcal{M}\triangleq\{1,2,...,M\}$, from which the drones depart from and return to, are located at fixed coordinates on the map.
Time periods and slots. We assume a set of time periods as $\mathcal{T}\triangleq\{1,2,...,T\}$. Each time period can be divided into a set of equal-length scheduling timeslots $\mathcal{S}\triangleq\{1,2,...,S\}$. In each timeslot, a drone can be controlled to fly to a cell and hover to collect sensor data.
Action. To explain the long-term navigation and sensing over all periods, we introduce the period-by-period actions of each drone $a^{u}=\{0,1,2,...,8\}$, $u\in\mathcal{U}$, meaning to control $u$ to move horizontally along eight directions, which are 1 = north (N), 2 = east (E), 3 = south (S), 4 = west (W), 5 = northeast (NE), 6 = southeast (SE), 7 = southwest (SW), and 8 = northwest (NW), or return to the origin ($a^{u}=0$). Under each action, $u$ executes a short-term navigation and sensing. After completing its sensing tasks, $u$ flies back to one of the charging stations to recharge fully and resume work in the next period $t+1$.
Matrix of plans. To explain the short-term navigation and sensing of drones over the cells and timeslots in a period $t$, $t\in\mathcal{T}$, the plan of a drone $u$ that travels through the direction of $a^{u}$ is introduced, denoted by $P^{u}(a^{u},t)$. It represents the specific navigation and sensing details, including the visited cells and corresponding energy consumption. The plan $P^{u}(a^{u},t)$ is encoded by a matrix of size $N\times S$, with each element represented as $p^{u}_{n,s}(a^{u},t)\in\{0,1\}$. Here, $p^{u}_{n,s}(a^{u},t)=1$ denotes that the drone $u$ hovers and collects all required data at cell $n$ at timeslot $s$, whereas for $p^{u}_{n,s}(a^{u},t)=0$ the drone does not hover at that cell at that time. Moreover, $P^{-u}(a^{u},t)=\{p^{-u}_{n,s}(a^{u},t)|_{n\in\mathcal{N},s\in\mathcal{S}}\}$ denotes the observed plan by drone $u$ in time period $t$, indicating that it incorporates and sums the plans of all other drones excluding its own. We formulate the aggregated plans of all drones observed by $u$ at cell $n$ and timeslot $s$ as follows:
$p_{n,s}(a^{u},t)=p^{u}_{n,s}(a^{u},t)+p^{-u}_{n,s}(a^{u},t).$ (1)
Matrix of required sensing values. In the context of a sensing task, each cell at a timeslot has specific sensing requirements that determine the data acquisition goal of drones. Such sensing requirements can be determined by city authorities as a continuous kernel density estimation, for example, monitoring cycling risk based on requirements calculated by past bike accident data and other information [15, 32]. The high risk level of a cell at a timeslot represents the high importance of sensing (e.g., the crucial intersection of traffic flow), and thus a high number of required sensing values is set. The matrix of required sensing values is denoted as $V(t)=\{V_{n,s}(t)|_{n\in\mathcal{N},s\in\mathcal{S}}\}$, where $V_{n,s}(t)$ denotes the required sensing value at cell $n$ and timeslot $s$ in the time period $t$. Based on actual sensing performance of drones according to Eq.(1), the sensing value collected by all drones observed by $u$ at cell $n$ and timeslot $s$ is formulated as:
$v_{n,s}(a^{u},t)=p_{n,s}(a^{u},t)\cdot V_{n,s}(t).$ (2)
Matrix of target. In a real-world scenario, drones lack prior knowledge of the amount of required sensing values before they begin their sensing operations. Therefore, it becomes essential to build a target on the fly that instructs the drones regarding when and where they should or should not fly to given information from the environment. The matrix of the target is defined as $\mathcal{R}(t)$, which denotes the sensing requirements for all drones at period $t$. The element of the target is denoted as $r_{n,s}(t)=\{0,1\}$, $n\in\mathcal{N}$, $s\in\mathcal{S}$, where $r_{n,s}(t)=1$ requires only a drone to visit the cell $n$ at timeslot $s$, and $r_{n,s}(t)=0$ does not require sensor data collection by a drone.
Assumptions. To simplify the scenario, this paper makes the following assumptions: (i) Each drone is programmed to fly at a distinct altitude during its movement between cells to prevent mid-air collisions risks [33]. While previous work [34, 35] has explored a safe and cost-effective approach to a more realistic collision avoidance, this is not the primary focus of this paper. (ii) Each time period concludes a flying period and a charging period: the time that drones perform sensing, and the time for charging. All drones finish charging before the next time period begins. We assume that each charging period is of equal duration and provides sufficient time for the drones to be fully charged. (iii) Each charging station is adequately equipped with charging capacity, enabling multiple drones to charge simultaneously without the need for queuing. This arrangement prevents any delays in the charging process.
III-B Problem Formulation
We model the scenario of a swarm of drones that perform sensing by considering the following performance metrics: 1) mission efficiency; 2) sensing accuracy; and 3) energy cost.
Mission efficiency. It denotes the ratio of sensing values in all cells at all timeslots collected by the drones during their mission over the total required values in all cells at all timeslots during the period $t$. The purpose is to collect sensing data as much as possible. It is formulated as:
$\mathsf{Eff}(a^{u},t)=\frac{\sum^{N}_{n=1}\sum^{S}_{s=1}v_{n,s}(a^{u},t)}{\sum^{N}_{n=1}\sum^{S}_{s=1}V_{n,s}(t)}.$ (3)
However, maximizing this metric leads to sensing imbalance and even blind areas. Some cells may be covered for a long time while some other cells may be never covered, i.e., the over-sensing and under-sensing [24]. Thus, it is necessary to consider matching indicators such as the sensing accuracy [24] to access such imbalances.
Sensing accuracy. It denotes the matching (correlation [^1] Error and correlation metrics (e.g. root mean squared error, cross-correlation or residuals of summed squares) estimate the matching, shown to be NP-hard combinatorial optimization problem in this context [7, 8].) between the total sensing values collected and the required ones. The metric is formulated as follows:
$\mathsf{Acc}(a^{u},t)=\sqrt{\frac{N\cdot S}{\sum^{N}_{n=1}\sum^{S}_{s=1}[v_{n,s}(a^{u},t)-V_{n,s}(t)]^{2}}}.$ (4)
Apart from improving the effciency and accuracy of drone sensing, the energy consumed by drones needs to be saved.
Energy cost. It is the battery usage of drones to perform spatio-temporal sensing. A power consumption model [36] is used to calculate the power consumption with input the specification of drones (weight, propeller and power efficiency). This model estimates the cost of navigation and sensing plans, and emulates the outdoor environments [15]. The energy cost of each drone $u$ is formulated as:
$E(a^{u},t)=\frac{C^{\mathsf{f}}(u)\cdot t^{\mathsf{f}}(a^{u})+C^{\mathsf{h}}(u)\cdot t^{\mathsf{h}}(a^{u})}{B(u)},$ (5)
where $B(u)$ is the battery capacity of drone $u$; $C^{\mathsf{f}}(u)$ and $C^{\mathsf{h}}(u)$ denote the flying and hovering power consumption of $u$ respectively; $t^{\mathsf{f}}$ and $t^{\mathsf{h}}$ are the flying and hovering time respectively, which are determined by the sensing plan $P^{u}(a^{u},t)$.
As described before, the optimization problem has three objectives: 1) maximize mission efficiency; 2) maximize sensing accuracy; and 3) minimize energy cost. Simultaneously achieving all of these objectives is a daunting task, as they often oppose each other. Maximizing both efficiency and accuracy is a trade-off. In addition, in order to achieve both of these goals, drones have to continuously move to collect data from different cells, which results in higher energy costs due to long-distance flights. However, some trips of drones may not improve sensing quality in the long run. For instance, at the end of a period, drones opt to return to the nearest charging station where there are fewer sensing tasks in the next period. This choice reduces the energy cost of the current period but comes at the significant expense of sacrificing the sensing quality in the next period. Therefore, the overall sensing performance is optimized by integrating these opposing objectives and selecting the plans of each drone.
Overall performance. The objective function is defined as the weighted sum of mission efficiency, sensing accuracy and energy cost:
$\begin{split}\mathop{\arg\max}\limits_{a^{u},u\in\mathcal{U}}\sum^{U}_{u=1}\sum^{T}_{t=1}[\;&\alpha_{1}\mathsf{Eff}(a^{u},t)+\alpha_{2}\mathsf{Acc}(a^{u},t)\\ &-\alpha_{3}E(a^{u},t)].\end{split}$ (6)
IV The Proposed Solution
Figure 1: System framework of DO-RL. The overall flow of DO-RL is depicted at the bottom; The plan selection is illustrated for two drones; The plan generation part outlines the procedure for generating a plan when drone $1$ is traveling to the east.
IV-A System Overview
In our problem, traditional DRL approaches suffer from the large action space and dimensional curse caused by the increasing number of time periods, leading to decline in sensing quality. To ensure high-quality sensing performance, the proposed DO-RL integrates decentralized drone coordination for short-term navigation and sensing optimization with long-term scheduling of flying directions. Fig.1 illustrates the designed system framework of DO-RL, consisting of three main components:
Plan generation. This component generates the navigation and sensing options for drones. Given the sensing map, each drone autonomously generates a finite number of discrete plans to initialize the overall process. This provides flexibility for the drones at the next stage to choose in a coordinated way. The generated plans are grouped into different flying directions.
Plan selection. This distributed component leverages a collective learning to coordinate drones to locally select the optimal navigation and sensing options from their generated plans within a period.
DRL-based scheduling. This component is the core of the framework, which leverages the DRL algorithm to enable drones to take the actions of overall flying directions between departure and destination charging stations in each time period. Thus, these actions are executed period-by-period, with each made only after the drone completes its sensing missions in the current period. The component includes two elements: a periodic state update, which updates the state of drones for DRL-based scheduling in the next period, and a multi-agent reinforcement learning module, which is built based on centralized training and decentralized execution.
The DO-RL process works as follows: Each drone initially generates a set of plans. It then selects a flying direction and chooses a subset of plans aligned with that direction. Each drone autonomously picks a plan from this subset, shares it with the swarm, and observes others’ choices. Based on its action and observations, the drone calculates a reward function and updates its state (including location, battery level, and sensing requirements), and stores these results in a multi-agent reinforcement learning buffer to refine its flight strategies. This cycle of DRL-based scheduling and plan selection repeats at each time period until the mission is complete. However, DO-RL requires frequent information sharing, which makes communication inefficient and vulnerable to potential failure of individual drones.
Figure 2: Overview of the structured tree communication model. During the bottom-up process, Drone 3 aggregates the plans of its children, i.e., drone 1 and drone 2, and sends them to its parent agent with its own plan. During the top-down process, each parent agent sends the total aggregated plans to its children such that all agents obtain the observed plan.
IV-B Communication Model
To overcome these barriers, DO-RL leverages a structured tree communication model. Each drone, controlled by a local agent, locates and connects with other agents into a tree communication structure, as shown in Fig. 2. Once agents have established their proximity-based relationship, e.g., Euclidean distance, they are positioned starting from the leaves up to the root, each interacting with its children and parent in a bottom-up and top-down [11, 10]. For example, a drone $u$ aggregates and shares its plan $P^{u}$ in a bottom-up fashion, and then obtains the observed plan $P^{-u}$ in the top-down process [7]. These plans serve as the observation of an agent in DRL modeling. As this coordination among agents using hierarchical organization reduces the need for every node to communicate with every other node, the system is highly-efficient in collaboration and information exchange [7].
Through iterative bottom-up and top-down exchanges, agents can flexibly change their plans and adapt to dynamic sensing environments. This ensures that even if some nodes fail, the overall sensing performance remains effective. Moreover, the system can dynamically reconfigure, repositioning drones within the tree in response to communication failures, ensuring that a single drone failure has minimal impact on information flow. The resilience of this model has been previously demonstrated in [11] with large-scale real-world datasets, and thus this paper focuses on sensing performance and scalability of the proposed solution.
V Detailed System Design
V-A DRL Modeling
The core design of DO-RL lies in applying DRL on our problem. Once drones take actions of flying directions, they execute plans and return to charging stations, changing their current locations and battery levels while observing the navigation and sensing of other drones. Therefore, the problem scenario can be explained into a Markov decision process [6]. We model the problem using state, action, and reward concepts: (i) State: The state $\textbf{s}_{t}$ at period $t$ consists of four components $(\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4})$, where $\mathcal{S}_{1}=\{c^{u}(t)|_{u\in\mathcal{U}}\}$ is the current locations of drones. Since drones charge at charging stations before taking actions, the location can serve as the index of the charging station. $\mathcal{S}_{2}=\{e^{u}(t)|_{u\in\mathcal{U}}\}$ is the current battery levels of drones, which are calculated based on the energy cost. $\mathcal{S}_{3}=\{P^{u}(t)|_{u\in\mathcal{U}}\}$ is the plan of $u$. $\mathcal{S}_{4}=\{P^{-u}(t)|_{u\in\mathcal{U}}\}$ is the aggregated plan of other drones excluding $u$, which are shared via the structured communication model. (ii) Action: The action $\textbf{a}_{t}=\{a^{1}_{t},...,a^{U}_{t}\}$ at period $t$ consists of period-by-period flying directions (N, E, S, W, NE, SE, SW, NW or O) determined by drones. (iii) Reward: Based on the objective function of Eq.(6), the expected immediate local reward of one drone at period $t$ is defined as follows: $\textbf{r}^{u}_{t}=\alpha_{1}\mathsf{Eff}(a^{u},t)+\alpha_{2}\cdot\mathsf{Acc}(a^{u},t)-\alpha_{3}E(a^{u},t)$. Throughout the training process or episodes, the overall reward fluctuates based on the actions taken by the drones. This helps drones in prioritizing areas rich in sensor data by maximizing their reward, i.e., the highest cumulative overall performance.
Figure 3: Process of path finding in plan generation.
V-B Plan Generation
The actions enable each drone to generate nine groups of plans, each associated with an action for a flying direction. These groups consist of multiple individual plans of the same size. A group of plans generated by drone $u$ during a period is defined as $G^{a^{u}}(t)$, where each plan in the group has the same action, $P^{u}(a^{u},t)\in G^{a^{u}}(t)$. Then, a plan can be defined as $P^{u}(a^{u},t)=\{l,K(a^{u}),J(a^{u}),E(a^{u},t)\}$, where $l$ is the index of the plan, $l\leq L$, ($L$ denotes the total number of plans in a group); $K(a^{u})$ denotes the indexes of cells within the searching range along flying direction $a^{u}$; $J(a^{u})$ denotes the indexes of visited cells within $K(a^{u})$; and $E(a^{u},t)$ denotes the energy consumed by $u$ traveling over $J(a^{u})$, i.e., the cost of the plan.
The details of generating a single plan are defined by the chosen path and the corresponding energy required to traverse that path, as shown in Fig. 3. As an example of a plan, the drone $u$ firstly finds a searching range to cells $K(a^{u})$: it determines a range with a certain width (equal to half of distance to nearest charging station) along its flying direction, and then searches the cells within this range. Next, it selects the visited cells $J(a^{u})$ randomly from this range. Then, $u$ finds the shortest possible path among $J(a^{u})$ via the Dijkstra’s algorithm for the Traveling Salesmen Problem. The path includes the information of which cell and which timeslot the drone uses to collect data, and thus both the traveling time $t^{\mathsf{f}}(a^{u})$ and sensing time $t^{\mathsf{h}}(a^{u})$ are calculated. The total energy consumption $E(a^{u},t)$ when $u$ travels over the path is then calculated via Eq.(5).
V-C Plan Selection
Given the plans in selected flying directions within a period, drones can adapt their assignments by selecting appropriate plans in response to changes in task targets, enhancing their adaptability in dynamic environments. In this process, the agents of drones improve their plan selections via tree communication. Specifically, each agent obtains the aggregated choices, i.e., the aggregated plan $P^{-u}(a^{u},t)$, from other agents, and chooses one of the plans $P^{u}(a^{u},t,l)$ such that all choices together $P^{u}(a^{u},t)+P^{-u}(a^{u},t)$ add up to match a target $\mathcal{R}(t)$. The target is used to steer each drone to sense over a unique cell during a timeslot, thereby avoiding the simultaneous sensing of multiple drones within the same cell, i.e., preventing over-sensing and under-sensing.
The purpose of coordination is to match the aggregated plans of all agents to the target while minimizing the energy cost $E(a^{u},t)$ of the plan selected by the drone. The cost function for each agent is formulated using the root mean square error (RMSE):
$\begin{split}\mathop{\min}\limits_{a^{u},u\in\mathcal{U}}\;&(1-\beta)\cdot\sqrt{\frac{\sum^{N}_{n=1}\sum^{S}_{s=1}[p_{n,s}(a^{u},t)-r_{n,s}(t)]^{2}}{N\cdot S}}\\ &+\beta\cdot E(a^{u},t),\end{split}$ (7)
where $\beta$ represents the behavior of an agent. As the value of $\beta$ increases, the agent becomes increasingly selfish, prioritizing plans with lower energy costs at the expense of higher root mean squared error. Each agent (or drone) aims to minimize the system-wide cost of sensing while considering its energy consumption. By minimizing this cost function, each drone coordinates to select its optimal plan, indicating a specific path within a period (comprising multiple timeslots). Via planning and selection, drones using DO-RL only need to take high-level actions of flying directions in each period, and thus reducing the number of decisions and observations in DRL.
V-D Periodic State Update
After choosing a plan, each drone changes its current state at the next period $t+1$, including its location $c^{u}(t+1)$, battery level $e^{u}(t+1)$, selected plan $P^{u}(t+1):=P^{u}(a^{u},t)$, and observed plan $P^{-u}(t+1):=P^{-u}(a^{u},t)$. However, due to dynamic changing environment, drones have no knowledge about the required sensing value in the next period. They need to predict and estimate the required sensing value to calculate reward function via Eq.(3) and (4). Depending on the predicted distribution of sensor data, the target is required to be updated.
Predicted sensing value. The predicted sensing value at period $t$, denoted as $\hat{V}(t)$, is updated in a time-reverse decay, formulated as follows:
$\hat{V}_{n,s}(t)=\sum^{t}_{t^{\prime}=1}(T-t+t^{\prime})\cdot\omega_{n,s}(t^{\prime})\cdot v^{\prime}_{n,s}(a^{u},t^{\prime}):=V_{n,s}(t),$ (8)
where $V^{\prime}_{n,s}$ denotes the data values collected by drones once they execute their selected plans; $\omega_{n,s}(t^{\prime})$ is a prediction coefficient such that $10<\omega_{n,s}(t^{\prime})<$ 1, $t^{\prime}\leq t$. To achieve accurate predictions, DO-RL leverages the Ordinary Least Squares regression method (OLS) to train these coefficients initially, and use the past experienced observations as the target distribution for training.
Target. The target $\mathcal{R}$ in the plan selection needs to be iteratively updated to steer the coordination of drones to choose plans for areas and timeslots with abundant sensor data. The percentile-based data filtering is used to eliminate the extreme low sensing values collected by drones. This approach effectively removes the need for data collection in regions with low sensing requirements (e.g., traffic exclusion zones). As a consequence, it helps drones in priortizing their operations in cells and time where sensing values are significantly high. We initialize the target as $r_{n,s}(0)=1$ to encompass all cells and timeslots, and update it as follows:
$\begin{split}r_{n,s}(t)=\left\{\begin{array}[]{ll}0,&v^{\prime}_{n,s}(a^{u},t)<\overline{v}\wedge p_{n,s}(a^{u},t)>0\\ r_{n,s}(t-1),&otherwise\end{array},\right.\end{split}$ (9)
where the threshold $\overline{v}$ is set iteratively and calculated as the value at the $100(1-U/N)$th percentile among the predicted data values $\hat{V}_{n,s}(t)$.
At the end of state update in DO-RL, the current collected data values $V^{\prime}_{n,s}$, the predicted values $\hat{V}_{n,s}$ and the target $\mathcal{R}$ are shared with each agent via the top-down interactions within the tree communication structure. This information will be stored and later sampled for multi-agent reinforcement learning.
V-E Multi-agent Reinforcement Learning
Our approach provides a generalized model for multi-agent reinforcement learning by adopting the framework of centralized training and decentralized execution, as shown in Fig.4. During execution, each agent obtains a local aggregated observation comprising both private and public information. Apart from their own private data such as locations, battery levels, and selected plans, this observation contains aggregated plans shared via the tree communication model (shown in Fig. 2). The training part is built based on an actor-critic model. The actors allow drones to work independently and choose actions based on their observations and policies, allowing for parallel data collection and processing. The critic, however, evaluates the actions of drones and guide them toward collectively optimal decisions. In addition, DO-RL employs Proximal Policy Optimization (PPO) to prevent detrimental updates and improving the stability of the learning process [37].
Figure 4: DRL-based long-term scheduling overview.
The process of training primarily involves the following steps: Firstly, each agent takes an action for a flying direction based on the aggregated observation $\textbf{o}^{u}_{t}$ shared via coordination as well as the reward $\textbf{r}^{u}_{t}$. Once all actions are determined, drones select their plans and transition to a new state. Subsequently, the buffer, a data storage structure used for experience replay, stores all transitions of each agent $(\textbf{o}^{u}_{t},\textbf{a}^{u}_{t},\textbf{r}^{u}_{t},\textbf{o}^{u}_{t+1})$. Several groups of transitions $(\textbf{o}^{u}_{k},\textbf{a}^{u}_{k},\textbf{r}^{u}_{k},\textbf{o}^{u}_{k+1})$ are sampled randomly ($H$ groups of transitions) for updating the parameters of both the critic and actor networks. The algorithm is typically an extension of the actor-critic policy gradient approach, employing two deep neural networks for each agent: a critic network $Q(\cdot)$ and an actor network $\pi(\cdot)$. The critic network estimates the reward associated with a transition using Bellman equation, and its parameter $\theta^{Q}$ is then updated by minimizing a loss function $L(\theta^{Q})$:
$L(\theta^{Q})=\frac{1}{H\cdot U}\sum^{H}_{k=1}\sum^{U}_{u=1}(\hat{A}^{u}_{k})^{2},$ (10)
$\hat{A}^{u}_{k}=\textbf{r}^{u}_{k}+\gamma\cdot Q(\textbf{o}^{u}_{k+1},a^{u}_{k+1})-Q(\textbf{o}^{u}_{k},a^{u}_{k}),$ (11)
where $\gamma$ is a discount factor; $\hat{A}^{u}_{k}$ is an advantage function. Then, the critic network provides $\hat{A}^{u}_{k}$ to the actor network to increase the probability of actions that have a positive impact and decrease the ones that have negative impact. The actions are taken by drones as $a^{u}_{t}=\pi(\textbf{o}^{u}_{t})$. The parameter $\theta^{\pi}$ of the actor network is updated by maximizing the clip objective $L_{CLIP}(\theta^{\pi})$, as follows [37]:
$\begin{split}L_{CLIP}(\theta^{\pi})=\;&\frac{1}{H\cdot U}\sum^{H}_{k=1}\sum^{U}_{u=1}min(ratio(\theta^{\pi},u)\hat{A}^{u}_{k},\\ &clip(ratio(\theta^{\pi},u),1-\epsilon,1+\epsilon)\hat{A}^{u}_{k}),\end{split}$ (12)
where $\epsilon$ is a hyperparameter; $clip(\cdot)$ defines the surrogate objective by limiting the range of $ratio(\theta^{\pi},u)=\frac{\pi(a^{u}_{k}|\textbf{o}^{u}_{k})}{\pi_{old}(a^{u}_{k}|\textbf{o}^{u}_{k})}$ using clipping, thereby eliminating incentives to exceed the interval $[1-\epsilon,1+\epsilon]$; $\pi_{old}$ denotes the older policy of the actor network in the previous iteration.
Figure 5: The central business district of Munich, Germany: (a) Basic scenario with $64$ cells, $4$ charging stations and high density of vehicles; (b) Increase the number of cells to 100; (c) Increase the number of charging stations to 9; (d) Change to a new map with low density of vehicles.
VI Performance Evaluation
In this section, an overview of the experimental settings is presented. The sensor dataset, specification of drones, optimization algorithm employed, and the used neural network are introduced. Then, the baselines and performance evaluation metrics are discussed. Finally, the results are assessed across various scenarios.
Figure 6: The distribution of both charging stations and traffic vehicles in the maps with high and low density of vehicles. There are $4$ charging stations uniformly distributed in the map with $64=8\times 8$ cells lined up over the map. The blue arrows symbolize the flow of vehicular traffic, with their length proportional to the volume of vehicles over $8$ periods.
VI-A Experimental Settings
Scenario model. In order to evaluate the DO-RL, we model a real-world transportation scenario, where a swarm of drones perform the sensing tasks of traffic monitoring. The number of vehicles serves as the required sensing value in the model. The scenario involves a real-world map of Munich city, imported to the simulations of urban mobility (SUMO) [^2] https://www.eclipse.org/sumo/ to accurately generate realistic flows of vehicles. Fig. 5 illustrates a selected map of $1600\times 1600$ meters in the city with the simulation time of $50$ hours, which is $100$ periods. It has a high density of vehicles, approximately $2,000$ vehicles passing by per hour. The area split into a finite number of cells, each is defined as a rectangular square with the size of $200\times 200$ meters that can be captured by the cameras of drones. Fig. 6 shows the distribution of both charging stations and traffic vehicles in the map. The charging stations are uniformly distributed in the map. We employ cross-validation: $80\%$ simulation time of the datasets for training and $20\%$ for testing. There are two scenarios for the experimental evaluation in this paper: (i) Basic scenario. It has $64$ cells, $4$ charging stations and $8$ time periods (set as $30min$ for each period). It has high density of vehicles, around $2,000$ vehicles passing by per time period. $16$ drones sense the area (camera recording) in parallel. (ii) Complex scenario. It varies the parameter settings such as the density of drones and vehicles, as well as the number of time periods, cells and charging stations. Noted that the simulated traffic flow can be displayed in a big screen in a testbed experimentation using hardware to emulate the complex outdoor sensing environments, which has been proved by our previous work in [15, 34].
Drones. We use DJI Phantom 4 Pro model [^3] https://www.dji.com/uk/phantom-4-pro/info. Each drone has a body weight of $1.07$ kg and four propellers, each with a diameter of $0.35$ m. The ground speed of the drones is set at $6.94$ m/s, and the drag force is determined to be $4.1134$[36]. As a consequence, the power consumption remains consistent across all drones. Additionally, drones are equipped with a $6000$ mAh LiPo 2S battery ($0.31$ kg). The maximum flying time for drones is approximately $30$ min, which is set as one flying period. A flying period has $S=30$ timeslots, each of one minute. Moreover, to ensure that the camera covers the entire area of a cell (see Fig. 5), the minimum hovering height of drones is determined at which the field of view of the camera and the cell overlap. The hovering height $h$ is calculated based on the distance between any two cells (approximately $200$ meters), the camera resolution, focal length derived from camera calibration, and ground sampling distance [38]. Therefore, each drone equipped with a 4K camera senses from a minimum height of $82.4$ meters.
Planning and distributed optimization. We leverage the decentralized multi-agent collective learning method of EPOS [^4] EPOS is open-source and available at: https://github.com/epournaras/EPOS., the Economic Planning and Optimized Selections [7]. We generate $L=64$ plans for each agent. All generated plans are made openly available to encourage further research on coordinated spatio-temporal sensing of drones [15, 24, 12]. Each agent in EPOS is mapped to a drone. During the coordinated plan selection via EPOS, agents self-organize into a balanced binary tree as a way of structuring efficient learning interactions [39]. The shared goal of the agents is to minimize the RMSE between the total sensing values collected and the target via Eq.(7). In the one execution of EPOS, the agents perform $40$ bottom-up and top-down learning iterations during which RMSE converges to a minimum optimized value.
Neural network and learning algorithm. In our problem, the weight parameters $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ are all set to a value of $1$. A total of $H=64$ groups of transitions are sampled as minibatches in a replay buffer, with a discount factor of $\gamma=0.95$ and a clip interval hyperparameter of $0.2$ for policy updating. We try the recurrent neural network (RNN), and use $W=64$ neurons in the three hidden layers of the RNN in both critic and actor networks. The activation function used for the networks is tanh. The models are trained over $\mathcal{E}=5000$ episodes, each consisting of multiple epochs (equal to the number of time periods).
Complexity. Given the number of nodes per layer $W$, number of episodes $\mathcal{E}$, size of state space $\mathcal{X}$, and number of actions $A$, the computational complexity of DNN is approximately $O(C_{dnn}):=O((\mathcal{X}+A)W+3W^{2})$ [25]. The comparison of both computational and communication cost is shown in Table II, where $I$ denotes the number of iterations in EPOS. The results illustrate that DO-RL significantly lowers the computation cost by reducing the number of actions required during training on a temporal scale, compared to the state-of-the-art MAPPO. DO-RL also has low communication cost due to the tree communication structure.
TABLE II: Comparison of computational and communication costs.
Attributes   Approaches.:
EPOS [7]
MAPPO [16]
DO-RL
Computational Cost $O(LI\log U)$ $O(\mathcal{E}TSC_{dnn})$ $O(\mathcal{E}T(C_{dnn}+LI\log U))$
Communication Cost $O(I\log U)$ $O(U^{2})$ $O(I\log U)$
Figure 7: High density of drones increases both mission efficiency and sensing accuracy of all methods, especially for short-term optimization methods (DO-RL and EPOS). DO-RL also shows its superior performance across different drones densities. Changing the drones density by increasing the number of drones from $8$ to $64$ and fixing 8 periods, 64 cells, 4 charging stations and high density of vehicles.
Figure 8: High number of periods increases both mission efficiency and sensing accuracy of all methods, especially for long-term learning methods (DO-RL and MAPPO). DO-RL also shows its superior performance across different number of periods. Changing the the number of periods from $4$ to $16$ and fixing 16 drones, 64 cells, 4 charging stations and high density of vehicles.
VI-B Baselines and Metrics
We compare DO-RL with the following baseline methods:
Greedy. It aims to assign each drone to a single cell, enabling it to accomplish the necessary sensing tasks within a time period [20]. This algorithm minimizes the number of visited cells and travel distance, resulting in reduced energy consumption for drones. To promote decentralization, we implement the algorithm with a local view for each drone, i.e., they have no coordination or awareness of cell occupancy by other drones. By adopting this approach, we ensure an energy-efficient and independent operation of each drone.
EPOS. It is designed to select the optimal plan for drones based on the generated plans [7]. It incorporates the plan generation and the periodic state update, but does not support any long-term strategic navigation. Agents randomly choose any charging station to land on at every period.
MAPPO. It is one state-of-the-art DRL algorithm using PPO [16, 37], but does not include distributed optimization compared to DO-RL. Moreover, agents in MAPPO learn the flying directions timeslot by timeslot, rather than period by period as in DO-RL. At each timeslot, a drone takes actions to horizontally move to an adjacent cell in eight directions or hover (similar to the actions in DO-RL), and returns to the nearest charging station at the end of every $S$ timeslots (one period). For fair comparison, this method employs the same reward function and structured tree communication model to share the aggregated observation.
The evaluation of all algorithms includes three key metrics: 1) Mission efficiency, 2) Sensing accuracy and 3) Energy cost. These metrics are determined using Eq.(3), (4) and (5) respectively. We use the number of vehicles detected/observed by drones to represent the sensing values defined in this paper. Furthermore, to obtain a comprehensive assessment that considers all three metrics, we conduct an overall performance evaluation using Eq.(6).
Figure 9: Both low number of cells and high number of charging stations increase the overall performance of all methods. Long-term learning methods outperform short-term optimization methods in observing traffic flow when vehicles are sparsely distributed across the map. Performance comparison under varying parameters: the number of cells ($64$ and $100$), the number of charging stations ($4$ and $9$), and the density of vehicles (high and low). The red palette on the right represents the values of energy cost, while the blue one denotes the values of mission efficiency, sensing mismatch and overall performance.
Figure 10: High number of drones/cells with fixed drones density increases the sensing accuracy of all methods, peaking at $36$ drones and $144$ cells. Changing the number of drones and cells but fixing the drones density at $0.25$ with 8 periods, 4 charging stations and high density of vehicles.
VI-C Results and Analysis
There are six dimensions for the complex scenario we study here: (i) the density of drones, (ii) the number of periods, (iii) the number of cells, (iv) the number of charging stations, (v) the density of vehicles, (vi) the number of drones and cells while fixing the density of drones, (vii) the remaining battery level, and (viii) the charging load. Results are shown in Fig. 7-12, where the shadow area around the lines represents the standard deviation error of the results.
Density of drones. It denotes the ratio of the number of drones over the number of cells, representing the coverage of a single drone over the map. In this case, we fix the number of cells as $64$, and increase the number of drones from $8$ to $64$, that is, the drones density increases from $0.125$ to $1.0$. Fig. 7 illustrates the performance of DO-RL and baseline methods when using different drones density to perform traffic monitoring. As drones density increases, both mission efficiency and sensing accuracy of DO-RL increase linearly. These are approximately $23.0\%$ and $15.8\%$ higher than EPOS respectively, as shown in Fig. 7 and Fig. 7  [^5] DO-RL, EPOS and MAPPO have similar overall performance when drones density is $1.0$, with maximum p-value of $0.04$, because there are ample drone resources available to effectively coordinate the coverage of the map.. This is because DO-RL controls the traveling direction of drones based on the predicted sensor data, so that drones cover the area with the maximum sensing value. MAPPO has high efficiency and accuracy with a low density of drones (maximum p-value less than $0.001$ using Mann-Whitney U test), but its performance degrades when the drones density is higher than $0.375$ due to the high computational complexity. In contrast, the plan selection in DO-RL reduces the action space and mitigates the learning difficulty, resulting in high performance with a high number of drones. In Fig. 7, the energy cost of DO-RL is on average $1.7\%$ lower than EPOS and $7.9\%$ lower than MAPPO. Greedy has the minimum energy cost among all methods, only $7.89\%$ lower than DO-RL, but has the lowest performance in efficiency and accuracy. DO-RL sacrifices a little energy for sensing, but still gets lower cost than other methods.
Number of periods. As shown in Fig. 8, DO-RL achieves superior performance compared to other methods, increasing linearly as the number of periods increase (increased by $51.3\%$ in quadruple time). Although it has statistically similar mission efficiency to MAPPO when the number of periods is $4$, this metric increases dramatically because drones update and obtain more accurate predicted sensor data with higher number of periods. The timeslot-by-timeslot learning in MAPPO perplexes the actions learned by drones compared to DO-RL, since drones only take actions once per period ($30min$) in DO-RL but take actions per minute in MAPPO. This also results in higher flying energy consumed by drones, and thus the energy cost of MAPPO is around $8.1\%$ higher than DO-RL.
Fig. 9 illustrates the performance comparison between DO-RL and baseline methods, varying the number of cells, charging stations and the density of vehicles, while keeping the number of drones fixed at $16$ and periods at $8$.
Number of cells. If the number of cells increases from $64$ to $100$, the density of drones decreases. In Fig. 9, DO-RL shows a lower decrease in mission efficiency (decreased by $27.04\%$) than MAPPO (decreased by $44.50\%$) and EPOS (decreased by $45.11\%$) as the number of cells increases. Furthermore, DO-RL keeps relatively constant both sensing accuracy and energy cost as the number of cells varies.
Number of charging stations. If the number of charging stations increases from $4$ to $9$, the distance between the sensing areas and stations is reduced, cutting down the energy cost of recharging. As a result, the overall performance of both DO-RL and EPOS increase by $19.52\%$ and $14.35\%$ respectively.
Density of vehicles. If we select a new map area that has low density of vehicles, the distribution of sensor data changes (with different road distribution), as shown in Fig. 6. In Fig. 9, both DO-RL and MAPPO have higher mission efficiency and sensing accuracy than the other two methods, with an average of $90.65\%$ higher efficiency and $76.85\%$ higher accuracy. This is because the learning methods observe the environment and control drones to collect data in cells and timeslots with the highest number of vehicles. DO-RL performs better than other methods under low vehicle density even though the number of cells and charging stations increase.
Number of drones/cells with fixed drones density. If we fix the density of drones but increase the number of drones, the number of cells increases proportionally. For example, we use the density of drones in the basic scenario of $0.25$, and increase the number of drones from $9$ to $49$ and the number of cells from $36$ to $196$. As shown in Fig. 10, when the number of drones/cells increases, the sensing accuracy of all methods increases. However, the mission efficiency of both EPOS and MAPPO decreases (by $35.93\%$ and $26.42\%$ respectively). This is due to EPOS facing challenges in efficiently scheduling drones for a large number of cells, while MAPPO experiences increased computational complexity. In contrast, DO-RL overcomes these challenges and demonstrates a $4.62\%$ increase in mission efficiency. When increasing to $49$ drones and $196$ cells, however, the sensing accuracy of DO-RL degrades, decreased by $15.96\%$, due to the high complexity of scheduling and computation. In overall, DO-RL has statistically higher overall performance than other methods, with an average improvement of $46.98\%$ (maximum p-value less than $0.001$).
(a) High density of vehicles.
(b) Low density of vehicles.
Figure 11: DO-RL keeps drones at a safe remaining battery level before recharging. Performance comparison of the remaining battery level of four methods on both high and low density of vehicles.
(a) Drones Density=0.25, high density of vehicles.
(b) Drones Density=0.25, low density of vehicles.
(c) Drones Density for DO-RL, high density of vehicles.
(d) Drones Density for DO-RL, low density of vehicles.
Figure 12: DO-RL relies on a lower number of charging places compared to other methods. Performance comparisons of the total energy demand of four methods with $0.25$ density of drones, see (a) and (b), and the DO-RL with different density of drones, see (c) and (d). Both comparisons are categorized by the high and low density of vehicles, see Fig. 6.
Remaining battery level. It denotes a percentage of battery level once drones have completed their sensing tasks. The value is calculated as an average among all drones within each time period. As shown in the Fig. 11(a) and 11(b), DO-RL keeps drones at a high remaining battery level on high and low density of vehicles, with approximately $30.25\%$ and only $1.05\%$ lower than Greedy. This level follows the drones’ safety regulations which suggest finishing the missions when battery life is around $25\%-30\%$. However, MAPPO does not meet the regulations under low density of vehicles, and the minimum battery level can reach $12.55\%$.
Charging load. It is the total energy demand of drones on each charging station over all time periods. This aims to study the placement of charging stations. As shown in the Fig. 12(a) and 12(b), the total energy demand on charging stations of DO-RL over all periods is more imbalanced compared to other methods. This is because with DO-RL drones learn to travel to the areas with the high required sensing data values. Since drones set the nearest charging stations as destinations, they rely on a lower number of charging places. For example, as shown in Fig. 12(c) and 12(d), when the density of drones is $0.125$ under high density of vehicles, drones only depart from and return to the charging station $D$ since the vehicle distribution around $D$ is more dense than other charging stations. As the density of drones becomes $1$, drones gradually rely on the charging station $B$ and $D$. Similarly, drones only need to fly over the vehicle-dense areas around the charging station $A$ and $B$ under low density of vehicles. These results provide insights to policy makers for providing higher amount of energy on vehicle-dense areas to support drones’ charging.
In summary, several scientific insights on experimental results are listed as follows: (i) Under scarce drone resources, it is more efficient and accurate to employ long-term learning methods for optimizing sensing. In contrast, under abundant drone resources, short-term optimization methods alone suffice. (ii) Long-term learning methods optimize drone resources usage for traffic management by directing drones towards regions with high vehicle density and advising on infrastructure planning for charging. (iii) Short-term optimization aids long-term learning by simplifying agent coordination, enhancing sensing quality, and preventing energy waste. This approach ensures that drones operate efficiently while maintaining battery levels above safety thresholds, leading to more reliable and effective missions.
VII Conclusion and Future Work
This paper studies a large-scale spatio-temporal task assignment problem by a swarm of autonomous drones. We propose a new synergetic optimization approach named DO-RL, by integrating both evolutionary distributed optimization and DRL, to solve the problem by optimizing the flight paths, data collection and recharging assignment of drone swarms. In each time period, drones determine their overall flying directions to destinations with charging stations via DRL, and autonomously evolve their navigation and sensing plans using collective learning. Extensive experimentation using realistic urban mobility reveals the effectiveness of DO-RL and provides valuable insights for the broader community: The combination of short-term and long-term strategies proves highly effective in overcoming their standalone limitations in drone-based traffic monitoring. Short-term methods compensate the challenges of training complexity, energy consumption and recharging, while long-term approaches account for maximizing accumulated rewards, enhancing the overall sensing performance. This collaborative approach allows the system to dynamically adapt to changing environments, efficiently managing varying drone resources and enabling extended operations.
However, the designed framework can be further improved towards several research avenues: 1) Use of simulated and real-world datasets in other applications, such as disaster response. 2) Consideration of more realistic factors such as obstacles and charging capability. 3) Research of the real-time drone testbed to validate the applicability of the proposed method.
ACKNOWLEDGEMENTS
This research is supported by a UKRI Future Leaders Fellowship (MR-/W009560/1): Digitally Assisted Collective Governance of Smart City Commons–ARTIO, and the European Union, under the Grant Agreement GA101081953 for the project H2OforAll—Innovative Integrated Tools and Technologies to Protect and Treat Drinking Water from Disinfection Byproducts (DBPs). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them. Funding for the work carried out by UK beneficiaries has been provided by UKRI under the UK government’s Horizon Europe funding guarantee [grant number 10043071]. Thanks to Manos Chaniotakis and Zeinab Nezami for the support on the transportation datasets.
Appendix A Mathematical Notations
Table III illustrates the list of mathematical notations used in this paper.
TABLE III: Notations.
Notation Explanation
$u,U,\mathcal{U}$ Index of a drone; total number of drones; set of drones
$m,M,\mathcal{M}$ Index of a charging station; total number of charging stations; set of charging stations
$n,N,\mathcal{N}$ Index of a grid cell; total number of grid cells; set of grid cells
$s,S,\mathcal{S}$ Index of a timeslot; total number of timeslots in a period; set of timeslots
$t,T,\mathcal{T}$ Index of a period; total number of periods; set of periods
$a^{u}$ The action or flying direction taken by $u$
$P^{u},p^{u}_{n,s}$ The plan of $u$; value of $P^{u}$ at cell $n$ and timeslot $s$
$P^{-u},p^{-u}_{n,s}$ The observed plan of drones by $u$ excluding $P^{u}$; value of $P^{-u}$ at cell $n$ and timeslot $s$
$p_{n,s}$ The aggregated plans of all drones at cell $n$ and timeslot $s$
$V_{n,s}$ Required sensing value at cell $n$ and timeslot $s$
$v_{n,s}$ Collected sensing value at cell $n$ and timeslot $s$
$\mathcal{R},r_{n,s}$ Target; target value at cell $n$ and timeslot $s$
$l,L$ Index of a plan in $G^{a^{u}}$; total number of plans in $G^{a^{u}}$
$C^{\mathsf{f}}(u),t^{\mathsf{f}}$ Flying power consumption of $u$; flying time
$C^{\mathsf{h}}(u),t^{\mathsf{h}}_{u}$ Hovering power consumption of $u$; hovering time
$B(u)$ Battery capacity of $u$
$\alpha_{1},\alpha_{2},\alpha_{3}$ Weight parameters of mission efficiency, sensing accuracy and energy cost in the main objective function respectively
$c^{u},e^{u}$ Current location of $u$; current energy consumption of $u$
$K(a^{u})$ Indexes of cells within the searching range along flying direction $a^{u}$
$J(a^{u})$ Indexes of visited cells within $K(a^{u})$
$\beta$ Behavior of an agent in planning optimization
$o^{u}_{t},r^{u}_{t}$ Observation of $u$ at period $t$; reward of $u$ at the period $t$
$k,H$ Index of a sampled transition; number of sampled transitions
$\gamma,\epsilon$ Discount factor; clip interval hyperparameter
$Q(\cdot),\theta^{Q}$ Critic network; parameter of critic network
$\pi(\cdot),\theta^{\pi}$ Actor network; parameter of actor network
$\omega_{n,s},\overline{v}$ Prediction coefficient to calculate $\hat{V}_{n,s}$; threshold to remove the regions with low sensing requirements
$I,\mathcal{X},A$ Total number of iterations in EPOS; size of state space; number of actions in DRL
$\mathcal{E},W$ Number of episodes; number of neutrals per layer in DNN
Appendix B Pseudocode of Algorithm
we present the pseudocode of the proposed DO-RL shown in Algorithm 1.
1 Initialize critic network $Q(\cdot)$, actor network $\pi(\cdot)$ with weights $\theta^{Q}$, $\theta^{\pi}$, and their two target networks $Q^{\prime}(\cdot)$, $\pi^{\prime}(\cdot)$ with parameters $\theta^{Q^{\prime}}:=\theta^{Q}$, $\theta^{\pi^{\prime}}:=\theta^{\pi}$;
2 for episode $:=1$ to max-episode-number do
3       Reset environment and obtain the initial state $\textbf{s}_{1}$;
4       for period $t:=1$ to episode-length do
5             for drone $u:=1,...,U$ do
6                   Take action $a^{u}_{t}=\pi(\textbf{o}^{u}_{t}|\theta^{\pi})$;
7                   Select the set of plans $G^{a^{u}}(t)$;
8                   Pick a plan $P^{u}(a^{u},t)$ and share it with other drones via tree communication;
9                   Observe the aggregated plan $P^{-u}(a^{u},t)$ and pick an optimized plan through collective learning [7];
10                   Compute reward $\textbf{r}^{u}_{t}=\alpha_{1}\mathsf{Eff}(a^{u},t)+\alpha_{2}\cdot\mathsf{Acc}(a^{u},t)-\alpha_{3}E(a^{u},t)$;
11                   Store transition sample $(\textbf{o}^{u}_{t},\textbf{a}^{u}_{t},\textbf{r}^{u}_{t},\textbf{o}^{u}_{t+1})$ into buffer;
12                   Sample a random minibatch of $H$ samples $(\textbf{o}^{u}_{k},\textbf{a}^{u}_{k},\textbf{r}^{u}_{k},\textbf{o}^{u}_{k+1})$ from buffer;
13
14             end for
15
16       end for
17      Update $\theta^{Q}$ by minimizing the loss function via Eq.(10);
18       Update $\theta^{\pi}$ by maximizing the clip objective via Eq.(12);
19       Update the target networks: $\theta^{Q^{\prime}}\leftarrow\theta^{Q}$, $\theta^{\pi^{\prime}}\leftarrow\theta^{\pi}$;
20
21 end for
Algorithm 1 DO-RL
Appendix C Effect of different parameters
The objective of this appendix is to analyze the impact of various parameters within the plan generation and distributed optimization of the proposed method. The insights derived from this appendix can be used to make informed empirical decisions regarding parameter selection for the evaluation scenarios. The calculations presented herein can be automated for any scenario during hyper-parameter optimization. Such optimization is not the focus of this paper.
We use the DO-RL and the basic scenario to test the mobility range and the agents’ behavior.
Mobility range. Fig. 13 illustrates the effect of different numbers of visited cells. The higher the number of cells a drone visits, the higher the flying energy is, and the higher the energy cost is. The high number of visited cells perplexes the spatio-temporal navigation and sensing of a drone, leading to over-sensing. Specifically, multiple drones visit the same cell simultaneously and waste energy on collecting the same data. In contrast, if drones visit only one cell, they are free from over-sensing, but have a high probability to miss the area with a high required sensing data due to its low mobility range. As a result, we empirically select $J(a^{u})=2$ for each drone to optimize the overall performance.
Agents’ behavior. Fig. 13 illustrates the effect of agents’ behavior by varying the parameter of $\beta$ [7]. As $\beta$ increases from 00 to $1$, agents reduce the energy cost of their selected plans, while decreasing both efficiency and accuracy. This is because, according to Eq. 7, drones with higher $\beta$ choose a plan with lower energy cost, which degrades the matching between the total sensing and the target. Since the overall performance for $\beta\in[0.1,0.8]$ is statistically similar, using the Mann-Whitney U test with average p-value of $0.0003$, drones are allowed to choose their behaviors within this range.
Figure 13: DO-RL has its peak overall performance when $J_{u}=2$, $\beta\in[0.1,0.8]$. Change the parameters of DO-RL in mobility range and agents’ behavior.
Appendix D Power consumption model
Drones spend energy to surpass gravity force and counter drag forces due to wind and forward motions. A drone controls the speed of each rotor to achieve the thrust $T$ and pitch $\theta$ necessary to stay aloft and travel forward at the desired velocity while balancing the weight and drag forces. For a drone with mass $m_{b}$ and its battery with mass $m_{c}$, we define the total required thrust as follows:
$\mathcal{T}=(m_{b}+m_{c})\cdot g+F_{d},$ (13)
where $g$ is the gravitational constant, and $F_{d}$ is the drag force that depends on air speed and air density. For steady flying, the drag force can be calculated by the pitch angle $\theta$ as:
$F_{d}=(m_{b}+m_{c})\cdot g\cdot tan(\theta).$ (14)
Based on the model in [36], the power consumption with forward velocity and forward pitch is given by:
$C^{\mathsf{f}}=(v\cdot sin\theta+v_{i})\cdot\frac{\mathcal{T}}{\epsilon},$ (15)
where $v$ is the average ground speed; $\epsilon$ is the overall power efficiency of the drone; $v_{i}$ is the induced velocity required for given $T$ and can be found by solving the nonlinear equation:
$v_{i}=\frac{2\cdot\mathcal{T}}{\pi\cdot d^{2}\cdot r\cdot\rho\cdot\sqrt{(v\cdot cos\theta)^{2}+(v\cdot sin\theta+v_{i})^{2}}},$ (16)
where $d$ and $r$ are the diameter and number of drone rotors; $\rho$ is the density of air.
Moreover, the power consumption for hovering of a drone is calculated by:
$C^{\mathsf{h}}=\frac{\mathcal{T}^{3/2}}{\epsilon\cdot\sqrt{\frac{1}{2}\pi\cdot d^{2}\cdot n\cdot\rho}}.$ (17)